OpenStudy (anonymous):

Simplify the expression √20d+4√125d-3√45d. A. 845d B. 11√5d C.(√20+4√125-3√45) √d D.13√5d

OpenStudy (anonymous):

Not C or D, re-worked the equation and got B

OpenStudy (campbell_st):

each radical can be simplified \[\sqrt{20d} = 2\sqrt{5d}...4\sqrt{125d} = 20\sqrt{5d}...3\sqrt{45d} = 9\sqrt{5d}\] hope this helps you get an answer